# Bitcoin Under Volatile Block Rewards: How Mempool Statistics Can Influence Bitcoin Mining

### v1

@@ 4. Simplified Volatile Reward Model
In this section, we provide a simplified approach to analyze mining strategies under the volatile reward model. We discuss how to implement a simplified environment that reflects the behavior of the Bitcoin mempool. Besides, we present a Markov Decision Process (MDP)-based tool that calculates a lower bound on the optimal time-averaged profit an adversary can achieve before a difficulty adjustment under the volatile reward model.
4.1. Simplified Mempool Implementation
@@
In an MDP implementation, when the agent, which represents the adversary in our implementation, selects an action, the environment transitions to a new state and provides a corresponding reward. Under a fixed reward model, the reward is determined by the number of adversarial and honest blocks mined. However, in a volatile reward model, the block reward is dependent not only on the number of blocks but also on the transaction fees included in those blocks. To reflect the effect of transaction fees on the block rewards, our implemented environment encompasses an equation that calculates the transaction fee included in a block as a function of block generation time. We refer to this equation as the time-fee equation. Deriving a time-fee equation compatible with the Bitcoin environment is challenging, as transaction fees vary across different periods due to various reasons such as fluctuations in Bitcoin’s supply and demand dynamics. This variability indicates that the optimal mining strategy for an adversary can differ significantly across different periods. To maximize its profit, the adversary must accurately predict the time-fee equation for the targeted attack period.
@@
For the simplified mempool implementation, we extracted historical data (Blo, 2024) on total transaction fees and block generation times for Bitcoin blocks mined across different periods. Using regression techniques, we derived the time-fee equation for the target period. This time-fee equation allows us to define the reward function for the MDP implementation. In Fig. 4, the total transaction fee of a block in BTC, denoted by $f(\cdot)$, is depicted as a function of its generation time in minutes, denoted by $t$, for two different periods. Each point represents a block, and the line represents the time-fee equation.
(a) October 8, 2023
@@
Fig. 4(a) corresponds to the Bitcoin blocks mined on October 8, 2023, which we denote as $\texttt{Day}_{1}$. The transaction fee per block during this day is relatively low. The corresponding linear regression line that represents the time-fee equation is $f_{1}(t)=0.0069t+0.0359$. As on average $1$ Bitcoin block is mined every $10$ minutes, the average transaction fee per block in this period is equal to $0.1049$ BTCs. Fig. 4(b) is related to the Bitcoin blocks mined on December 17, 2023, which we denote as $\texttt{Day}_{2}$. During this day, Bitcoin has experienced one of the periods with the highest transaction fee per block. The corresponding time-fee equation for this day is $f_{2}(t)=0.0659t+3.0457$, indicating that the average transaction fee per block is equal to $3.7052$ BTCs.
In this section, we modeled the Bitcoin mempool using linear regression, overlooking the concavity of the time-fee function, and employing only a single predictor of block generation time. To enhance mempool implementation, one can employ more advanced regression techniques as discussed in Appendix D.
@@ 4.2. MDP-Based Tool
In this section, we introduce an MDP-based tool to analyze the profitability of selfish mining before a difficulty adjustment under a highly simplified volatile block reward model. Our MDP-based tool is presented in (our, 2024). MDP can be used to determine the optimal policy when the number of possible states in the environment is limited to fewer than approximately $10^{6}$. Therefore, analyzing the volatile reward model with MDP requires simplifying the environment’s representation to fit within a limited number of states. Our MDP tool is designed to calculate a lower bound for selfish mining profitability under the volatile reward model, allowing us to determine the minimum amount of mining power required for a miner to be incentivized to deviate from the honest strategy.
In our simplified model, we consider two forks: an honest fork and an adversarial fork. The honest fork is publicly known to all miners, while the adversarial fork is known only to the selfish miner, who can strategically decide when to publish the blocks in the adversarial fork. To account for the volatile reward model, we classify the blocks in the adversarial fork into three categories:
  • Blocks with normal block generation time: These adversarial blocks have an average generation time of 10 minutes, as they are mined after another adversarial block, without any intervening honest blocks.
  • Blocks with extended block generation time: These adversarial blocks have longer generation times, as there is one honest block between them and their preceding adversarial blocks in the adversarial fork. On average, these blocks have a block generation time of 20 minutes.
  • Blocks with highly extended block generation time: These adversarial blocks have even longer generation times, with at least 2 honest blocks intervening between them and their preceding adversarial blocks in the adversarial fork. On average, these blocks have a block generation time of 30 minutes or more.
Blocks with extended or highly extended generation times offer higher fee rewards compared to those with normal times. Thus, to achieve the optimal mining strategy in this simplified model, the adversary should aim to maximize the number of extended and highly extended blocks. The reason we consider only up to blocks with a 30-minute generation time is to limit the state space.
@@
The actions available to the adversary are similar to those introduced in (Sapirshtein et al., 2017) and include: override, adopt, match, and wait. In our simplified environment, each state is represented by five elements: $(l_{\mathcal{A}},l_{\mathcal{H}},G,\texttt{Latest},\texttt{Match})$. Here, $l_{\mathcal{A}}$ and $l_{\mathcal{H}}$ denote the lengths of the adversarial and honest forks, respectively, and can range from 0 to maxForkLen, where maxForkLen represents the maximum fork length in this environment. If the adversarial fork reaches maxForkLen, the selfish miner should publish the fork; if the honest fork reaches this length, the adversary should adopt the honest fork. The parameter maxForkLen limits the state space to a finite number of states. $G$ is an array of length maxForkLen that stores information about the block generation times in the adversarial fork. Each element in $G$ can be 0, 1, or 2: if the $i^{\text{th}}$ element is 0, the corresponding block in the adversarial fork has a normal generation time; if the $i^{\text{th}}$ element is 1, the block has an extended generation time; and if the $i^{\text{th}}$ element is 2, the block has a highly extended generation time. Latest is an element indicating how many honest blocks have been mined since the last adversarial block. It can take values 0, 1, or 2, denoting that the next adversarial block has a normal, extended, or highly extended generation time, respectively. Lastly, Match is a boolean element indicating whether the match action is currently in progress.
4.3. Results and Implications: Selfish Mining Under a Simplified Volatile Model
To analyze selfish mining profitability over different periods, we first define base fee and fee increase rate. The base fee is the intercept of a linear time-fee equation that represents the average amount of remaining transaction fees available in the transaction pool immediately after a block is mined. The fee increase rate is the slope of a linear time-fee equation that represents the increase in the collected transaction fee of a block per unit of time. At $t=0$, the transaction reward of the next potential block is equal to the base fee. This reflects reality, as there are always some low-value transactions in the mempool that miners can use to fill their blocks. As block generation time increases, the collected transaction fee rises by the fee increase rate for each additional unit of time due to the arrival of new transactions.
The impact of the base fee on selfish mining profitability is similar to that of the block protocol reward, as a block always receives the base fee in addition to the protocol reward, regardless of the block generation time. The main factor that results in selfish mining profitability before difficulty adjustment under the volatile reward model is the fee increase rate. The higher the fee increase rate, the more profitable selfish mining would be. The intuition, as discussed in Section 3.3, is that selfish mining increases the block generation time before a difficulty adjustment mechanism. This increase in the block generation time can raise the collected fee of the adversarial block proportional to the fee increase rate. To intuitively analyze the effect of the time-fee equation on selfish mining profitability before the difficulty adjustment, we define the term orphaning fee ratio. This term is defined as the ratio of the average reward received by a block with a generation time of 20 minutes to the average reward received by a block with a generation time of 10 minutes. The former block represents an adversarial block that has managed to orphan its previous block, while the latter represents a block mined following the honest strategy. The greater the orphaning fee ratio exceeds 1, the more profitable selfish mining becomes. If the orphaning fee ratio is close to 1, the block reward does not increase substantially over time, implying that selfish mining may not be significantly profitable.
We compare selfish mining profits before difficulty adjustment during $\texttt{Day}_{1}$ and $\texttt{Day}_{2}$ introduced in Section 4.1, under two conditions: when the protocol reward is still substantial and after it has converged to zero. We normalize the time-averaged profit of the entire network to 1, indicating that if all miners mine honestly, the time-averaged profit for each miner will equal their mining share. In Fig. 5, the lower bound of the time-averaged profit that a selfish miner can achieve before the difficulty adjustment on $\texttt{Day}_{1}$ and $\texttt{Day}_{2}$, under both protocol rewards of 0 and 3.125 BTC, is depicted as a function of selfish miner’s mining share. This graph corresponds to the selfish miner with a communication capability equal to $1$. We proceed by analyzing the results depicted in Fig. 5. Let $R(t)$ denote the total reward of a block as a function of block generation time and $r^{\texttt{orphan}}$ denote the orphaning fee ratio. Assuming a protocol reward of $3.125$ BTCs per block, the block reward-time equation during $\texttt{Day}_{1}$ and $\texttt{Day}_{2}$ can be obtained as $R_{1}(t)=0.0069t+3.1609$ and $R_{2}(t)=0.0659t+6.1707$, respectively. In $\texttt{Day}_{1}$, where the transaction fee is negligible compared to the protocol reward, the orphaning fee ratio is $r_{1}^{\texttt{orphan}}=1.0213$, making selfish mining not substantially profitable before the difficulty adjustment. However, on $\texttt{Day}_{2}$, where the transaction fee is comparable to the block reward, the orphaning fee ratio $r_{2}^{\texttt{orphan}}=1.0965$. As $r_{2}^{\texttt{orphan}}>r_{1}^{\texttt{orphan}}$, under the assumption of $3.125$ BTCs protocol reward per block, selfish mining is more profitable before the difficulty adjustment on $\texttt{Day}_{2}$ compared to $\texttt{Day}_{1}$.
Figure 5. A lower bound for the time-averaged profit before difficulty adjustment.
Considering the scenario that the protocol reward has converged to 00, the block reward-time equations would be the same as the time-fee equations. Under this scenario, the orphaning fee ratios are $r_{1}^{\texttt{orphan}}=1.6578$ for $\texttt{Day}_{1}$ and $r_{2}^{\texttt{orphan}}=1.1779$ for $\texttt{Day}_{2}$, respectively. Although the transaction fees on $\texttt{Day}_{1}$ are much lower than transaction fees on $\texttt{Day}_{2}$, as $\texttt{Day}_{1}$ has a higher orphaning fee ratio, miners are more incentivized to conduct selfish mining on $\texttt{Day}_{1}$ once the protocol reward converges to 00.
These observations conclude that selfish mining before a difficulty adjustment becomes profitable on days when the ratio of the fee increase rate to the base reward (base fee plus protocol reward) is substantial. To meet this condition with a significant protocol reward, transaction fees must also be high and comparable to the protocol reward. However, if the protocol reward becomes negligible, the Bitcoin network becomes more vulnerable to selfish mining on days when the rate of increase in transaction fees is high, even if those days do not coincide with high transaction fees per block.

### v2

@@ 4. Simplified Volatile Reward Model
The WeRLman environment analyzes adversarial strategies by considering only two types of transactions: normal and whale transactions. In this section, we introduce a new volatile reward model, referred to as the simplified environment. This environment enhances the WeRLman environment by including a greater number of transaction types, each with different fee levels, to provide a more accurate simulation of the real-world mempool. Besides, we present a Markov Decision Process (MDP)-based tool that calculates a the optimal strategy under the simplified environment.
In the WeRLman model, the volatility of block rewards arises from the inclusion of whale transactions. However, it is important to emphasize that the primary source of volatility in block rewards during the transaction fee era is not the occasional arrival of whale transactions that offer significantly higher fees compared to others. Instead, the main factor driving this volatility is the competition among transactions to be included in a block as quickly as possible. This competition forces transactions to adjust their fees based on the state of the mempool. Over time, as more valuable transactions accumulate in the mempool, new transactions need to offer higher fees to have a chance of inclusion in the next block. Therefore, during periods of high demand, block rewards can rise significantly due to the presence of higher-fee transactions in the mempool.
An important factor that increases transaction inclusion demand is extended periods during which no blocks are generated. Due to the random nature of Bitcoin’s mining process, block generation times can vary widely, ranging from a few seconds to over an hour. When block generation times are longer, miners have more opportunities to include higher-fee transactions in their blocks, thus increasing block rewards. This phenomenon is evident in the Bitcoin blockchain, where blocks with significantly higher transaction fee rewards compared to their surrounding blocks often correspond to those with unusually long block generation times. In our simplified model, instead of focusing on the concept of whale transactions, we concentrate on block generation time to analyze Bitcoin mining strategies.
4.1. Simplified Mempool Environment
@@
In a volatile reward model, miners’ profits depend not only on the number of blocks they mine but also on the transaction fees included in those blocks. To reflect the effect of transaction fees on the block rewards, our simplified environment encompasses an equation that calculates the transaction fee included in a block as a function of block generation time. We refer to this equation as the time-fee equation. Deriving a time-fee equation compatible with the Bitcoin environment is challenging, as transaction fees vary across different periods due to various reasons such as fluctuations in Bitcoin’s supply and demand dynamics. This variability indicates that the optimal mining strategy for an adversary can differ significantly across different periods. To maximize its profit, the adversary must accurately predict the time-fee equation for the targeted attack period.
@@
For the simplified mempool implementation, we extracted historical data (Blo, 2024) on total transaction fees and block generation times for Bitcoin blocks mined across different periods. Using regression techniques, we derived the time-fee equation for the target period. This time-fee equation enables us to estimate the reward function and analyze the profitability of mining strategies. As an example, in Fig. 4, the total transaction fee of a block in BTC, denoted by fee, is depicted as a function of its generation time in minutes, denoted by $t$, for two different periods. Each point represents a block, and the line represents the time-fee equation. Fig. 4(a) corresponds to Bitcoin blocks mined on January 1, 2023, with relatively low transaction fees per block, while Fig. 4(b) corresponds to blocks mined on December 17, 2023, during a period of high transaction fees.
(a) January 1, 2023
@@
A linear time-fee equation can be expressed as $\texttt{fee}(t)=\texttt{fee}_{0}+r_{\texttt{fee}}t$, where $\texttt{fee}_{0}$ and $r_{\texttt{fee}}$ represent the base fee and fee increase rate, respectively. The base fee is the intercept of a linear time-fee equation that represents the average amount of remaining transaction fees available in the transaction pool immediately after a block is mined. The fee increase rate is the slope of a linear time-fee equation that represents the increase in the collected transaction fee of a block per unit of time. At $t=0$, the transaction reward of the next potential block is equal to the base fee. This reflects reality, as there are always some low-value transactions in the mempool that miners can use to fill their blocks. As block generation time increases, the collected transaction fee rises by the fee increase rate for each additional unit of time due to the arrival of new transactions.
In the simplified model, we use an MDP-based tool to analyze selfish mining profitability. MDP can be used to determine the optimal policy when the number of possible states in the environment is limited to fewer than approximately $10^{7}$. Therefore, analyzing the volatile reward model with MDP requires simplifying the environment’s time-fee equation to fit within a limited number of states. To this end, we assume that the time-fee equation is a linear function, and time is divided into $M\geq 2$ discrete steps of the form $\{t_{i}=i\Delta\mid i\in[M]\}$, $\Delta$ denotes the length of each time step. For a block with block generation time $t$, the transaction fee reward can be obtained as follows:
(3) $\texttt{fee}(t)=\begin{cases}\texttt{fee}(i\Delta),&\text{if }\enspace i\Delta\leq t<(i+1)\Delta\enspace\text{for}\enspace i\in[M-1]\\ \texttt{fee}((M-1)\Delta),&\text{if }\enspace(M-1)\Delta\leq t\end{cases}$
In our environment, block generation time follows the exponential distribution with rate $\lambda$.
@@ 4.2. MDP-Based Tool
In this section, we introduce an MDP-based tool to analyze the profitability of selfish mining under the simplified environemnt introduced in Section 4.1. Our MDP-based tool, presented in (our, 2024), is designed to calculate a lower bound for selfish mining profitability under the volatile reward model, enabling us to determine the security threshold both before and after the difficulty adjustment.
@@
In our simplified model, we consider two forks: an honest fork and an adversarial fork. The honest fork is publicly known to all miners, while the adversarial fork is known only to the selfish miner, who can strategically decide when to publish the blocks in the adversarial fork. The actions available to the adversary are similar to those introduced in (Sapirshtein et al., 2017) and include: override, adopt, match, and wait. Each state in our simplified environment is represented by five elements: $(l_{\mathcal{A}},l_{\mathcal{H}},T_{\texttt{total}},T_{\texttt{last}},\texttt{fork})$. Here, $l_{\mathcal{A}}$ and $l_{\mathcal{H}}$ denote the lengths of the adversarial and honest forks, respectively, and can range from 0 to maxForkLen, where maxForkLen represents the maximum fork length in this environment. If one of the fork reaches maxForkLen, the selfish miner must either publish its private fork or adopt the honest fork. $T_{\texttt{total}}$ represents the total sum of block generation time steps for all the blocks in the adversarial fork. Given the assumption of a linear time-fee equation, the reward of the adversarial fork can be calculated using $T_{\texttt{total}}$ without the need to store the block generation time for each block separately, as follows: $l_{\mathcal{A}}(R+\texttt{fee}_{0})+r_{\texttt{rate}}\Delta T_{\texttt{total}},$ where $R$ is the protocol reward per block. $T_{\texttt{last}}$ represents the time steps elapsed since the mining of the last adversarial block and is used to update the value of $T_{\texttt{total}}$ at each event of adversarial block generation. fork can take values 0, 1, or 2, indicating that the latest mined block is adversarial, the latest block is honest, or the action match is currently in progress, respectively.
4.3. Difficulty-Adjusted Selfish Mining in the Simplified Environment
[…]
(4) $\texttt{Profit}^{\mathcal{A}}(\pi)=\lim_{T\to\infty}{\frac{\sum_{t=1}^{T}{R_{\mathcal{A}}(t;\pi)}}{\sum_{t=1}^{T}{N_{\mathcal{A}}(t;\pi)+N_{\mathcal{H}}(t;\pi)}}}\enspace.$
Note that although the average canonical block generation time under difficulty-adjusted selfish mining remains equal to $\mathbf{t}^{B}_{\texttt{ideal}}$, the average total block generation time (considering both orphan and canonical blocks) is affected by the adversary’s strategy. Therefore, we cannot assume that the total block generation time follows the initial rate $\lambda$. Ensuring the correct block generation time is crucial, as it impacts the transaction fee reward. To maintain a constant transaction arrival rate, we adopt the same heuristic as the WeRLman paper. At each step of mining a new block $B$, we sample an exponential random variable with mean $\lambda$ to determine the block generation time of $B$, only if $B$ is a block with a new height [^7] The block height is the distance of the block from the genesis block. in the system. However, if there already exists a block in the opposite fork with the same height as $B$, the generation time is set to 00. This approach ensures the transaction arrival rate remains consistent regardless of the adversary’s strategy.
Security threshold (comparison with WeRLman model): We provide a comparison between the security threshold in our simplified environment and the non-predictable WeRLman environment. The reason our results should not be compared with the security threshold of the original WeRLman environment is that, unlike the original environment, our simplified environment does not grant the adversary predictive capabilities. These predictive capabilities could be incorporated into our model by allowing the adversary to know the block generation time of the next block; however, this would not reflect real-world practices.
To ensure a fair comparison, we adjust the parameters in our model so that they align with the WeRLman model. Let $F$ and $p_{\texttt{whale}}$ denote the additional fee value of a whale transaction and the probability of a whale transaction in the WeRLman model, respectively. Also, let $M$ denote the number of discrete time points in our model, $t_{M-1}$ the greatest time point in our model, and $\texttt{fee}(t)$ the time-fee equation. We set $t_{M-1}$ such that the probability of the block generation time being equal to or greater than $t_{M-1}$ matches $p_{\texttt{whale}}$. Specifically, $t_{M-1}=\frac{-\texttt{Ln}(p_{\texttt{whale}})}{\lambda}$. This ensures that the probability of mining a block with the highest possible block generation time (the block with the highest fee) in our model matches the probability of mining a whale-included block in the WeRLman model. Additionally, we configure the time-fee equation $\texttt{fee}(t)$ to satisfy $\frac{\texttt{fee}(t_{M-1})}{\texttt{fee}(t_{\texttt{average}})}=1+F$, where $t_{\texttt{average}}$ is the average block generation time according to the discrete time division in our model. This ensures that, in both models, the ratio of the fee of the most valuable block to the fee of a normal block is equal to $1+F$.
By setting the number of discrete time points $M=2$, our model exactly matches the non-predictable WeRLman model, as both models only include blocks with two distinct fee levels. Increasing the number of discrete time points allows our model to represent a wider variety of blocks with different rewards, enabling a more precise analysis of profitability in the volatile block reward setting. In Figure 5, we compare the security thresholds in the non-predictable WeRLman model and our simplified model, assuming a protocol reward of 00 and a probability of mining the most valuable block of $p=0.001$. To obtain the security threshold for our model, we used the MDP-based tool introduced in Section 4.2, with the number of time points set to $M=30$. As shown in Figure 5, considering an environment with higher block reward volatility decreases the security threshold. This occurs because, in a more volatile model, blocks with rewards lower than a normal 10-minute block exist, allowing the adversary to take the risk of withholding these blocks to orphan valuable honest blocks. Additionally, a greater number of blocks with rewards exceeding the normal reward exist, resulting in more frequent attractive opportunities to deviate from the honest strategy.
Figure 5. Security threshold comparison under our simplified model and non-predictable WeRLman model.
5. Pre-DAM Selfish Mining Profitability Under Volatile Reward Model
In this section, we demonstrate that selfish mining before difficulty adjustment can be profitable under the volatile reward model, enabling miners to deviate from the honest strategy for immediate gain. We first discuss the intuition of how selfish mining, whish is unprofitable under the fixed reward model, becomes immediately profitable and thus more threatening in the volatile reward model.
5.1. Selfish Mining Profit Lag Under the Fixed Block Reward Model
Since Bitcoin’s introduction, it has progressed without experiencing long-range or detectable selfish mining attacks. We argue that, in addition to the high hash power required for a successful attack, another contributing factor is the prolonged time required for selfish mining to become profitable under the fixed block reward model. As discussed in Section 2.3, selfish mining during the first difficulty epoch incurs financial losses for the adversary and must be sustained for several days to become profitable. Figure 6 shows the minimum number of days required for a selfish mining attack to exceed the profitability of honest mining under the fixed block reward model, a period referred to as the profit lag in (Grunspan and Pérez-Marco, 2023). The figure considers adversaries with communication capabilities of $\gamma=0.5$, $0.75$, and $1$ who follow the optimal selfish mining strategy from (Sapirshtein et al., 2017), which maximizes block ratio. The main takeaway is that, under the fixed block reward model, selfish mining demands a significant time investment (more than one ideal epoch or two weeks) to become profitable.
Under the volatile block reward model, selfish mining can yield immediate profits by shortening or eliminating the initial loss period, making it a more significant threat to the Bitcoin network. The primary reason for the immediate profitability of selfish mining under the volatile model is that the attack increases the average generation time of adversarial blocks before the first DAM. While this does not affect the protocol reward per block, it boosts the average transaction fee reward per adversarial block.
Let $\lambda_{1}$ [^8] Note that $\lambda_{1}=\lambda\alpha$, where $\lambda$ is the total block generation rate. and $R_{1}$ denote the adversarial canonical block mining rate and the average block reward under the volatile reward model when all the miners including the adversary mine honestly. Also, let $\lambda_{2}$ and $R_{2}$ denote the adversarial canonical block mining rate and the average block reward under the volatile reward model during the first epoch of selfish mining. As discussed in Section 2.3, the adversary’s time-averaged profit during the honest mining and the first epoch of selfish mining are equal to $\lambda_{1}R_{1}$ and $\lambda_{2}R_{2}$, respectively. By conducting a selfish mining attack, the adversarial canonical block generation rate decreases in the first epoch of the attack as some of the adversarial blocks may become orphaned, implying that $\lambda_{2}\leq\lambda_{1}$. However, unlike the fixed block reward model, $R_{2}$ is not necessarily equal to $R_{1}$. In the volatile model, as a result of the increase in the block generation time [^9] The network block generation time is inversely proportional to the total block generation rate. during the first epoch of the attack, the average block reward may increase, implying that $R_{2}\geq R_{1}$. The reason behind this increase in the block reward is intuitively illustrated in Fig. 7. Therefore, one of the time-averaged profits $\lambda_{i}R_{i}$ for $i\in\{1,2\}$ may be greater than the other depending on the adversary’s mining parameters and the mempool statistics. This shows that under the volatile reward block model, selfish mining can be profitable even in the first epoch of the attack.
5.3. Pre-DAM Selfish Mining in the Simplified Environment
In this section, we analyze the profitability of selfish mining before the difficulty adjustment under the simplified model introduced in Section 4 using our MDP-based tool. Let $\mathbf{t}^{B}_{\texttt{ideal}}$ denote Bitcoin’s ideal average block generation time, and $R_{\mathcal{A}}(t;\pi)$ represent the total reward included in adversarial blocks added to the canonical chain at time step $t$ under strategy $\pi$. Under pre-DAM selfish mining, the average block generation time (considering both orphan and canonical blocks) remains $\mathbf{t}^{B}_{\texttt{ideal}}$. Since a new block is mined at each step in the environment, the expected elapsed time up to step $T$ is $T\mathbf{t}^{B}_{\texttt{ideal}}$. By normalizing $\mathbf{t}^{B}_{\texttt{ideal}}$ to 1 and ignoring mining costs, the time-averaged profit of selfish mining before difficulty adjustment, i.e., the objective function, can be expressed as:
(5) $\texttt{Profit}^{\mathcal{A}}(\pi)=\lim_{T\to\infty}{\frac{\sum_{t=1}^{T}{R_{\mathcal{A}}(t;\pi)}}{T}}\enspace.$
Since, before DAM, the difficulty is not adjusted based on the adversary’s strategy, the generation time of each block follows an exponential distribution with rate $\lambda$. Therefore, at each step of mining a new block $B$, we sample a random number from an exponential distribution with rate $\lambda$ to represent its generation time.
Security threshold (upcoming years): To analyze the security threshold of Bitcoin in upcoming years, we need an estimation of the time-reward equation of Bitcoin blocks. The reward-fee equation takes the form $\texttt{Reward}(t)=R+\texttt{fee}_{0}+r_{\texttt{fee}}t,$ where $R$ is the protocol reward, $t$ is the block generation time in minutes, $\texttt{fee}_{0}$ is the base fee, and $r_{\texttt{fee}}$ is the fee increase rate. We extracted the Bitcoin block transaction fee for all blocks mined during the period from January 2023 to December 2024. Then, for each day in this period, we used regression techniques to obtain the reward-fee equation based on different values of the protocol reward $R$. For each protocol reward $R$, we calculated the average value of the block reward increase ratio defined as $\frac{r_{\texttt{fee}}}{R+\texttt{fee}_{0}}$ among those promising days whose block reward increase ratio is in the top 10 percent. According to our analysis, the block reward increase ratio during promising days under the current protocol reward of 3.125 BTC, after 1, 2, 3, 4 protocol reward halvings, and under the eventual protocol reward of 0, can be estimated as $0.0091$, $0.0150$, $0.0230$, $0.0336$, $0.0459$, and $0.2057$, respectively. Figure 1 illustrates the security threshold for Bitcoin mining before difficulty adjustment based on the extracted block reward increase ratios, considering three different adversarial communication capabilities of $0.5$, $0.75$, and $0.9$. Note that for the same block reward increase ratio, the security threshold for pre-DAM selfish mining is higher than that for the difficulty-adjusted case. For miners whose mining share exceeds the former threshold, selfish mining does not incur an initial loss period and becomes profitable immediately. However, for miners whose mining share falls between the two thresholds, selfish mining results in losses initially and becomes profitable only in the long term. We can conclude that in the coming years, as transaction fees gradually become the primary source of rewards, the threat of selfish mining becomes more significant. This is because not only does the minimum mining power required for a successful attack decrease, but the attacker would also gain immediate profits upon executing the attack.
[…]
Figure 8. Time-averaged profit before difficulty adjustment.